What is the value of x in a rightv triangle when length =6 and width=9
Answer by tutorcecilia(2152)

What is the value of x in a rightv triangle when length =6 and width=9
I am going to assume that the hypotenuse is "x".
Using the Pythagorean Theorem: a^+b^2=c^2
Insert the values: 6^2+9^2=c^2
Solving: 36+81=c^2
117=c^2
Square both sides:10.82=c
Plug-in and check:
6^2+9^2=10.82^2
36+81=117
36+81=117
117=117